OpenStudy (anonymous):

How do i solve x^4 - 2x^3 + 8x^2 - 32x - 128 = 0 in the real number system?

OpenStudy (anonymous):

x^4 - 2x^3 + 8x^2 - 32x - 128 = 0 By synthetic division 4 |.1 -2.. 8 -32 -128 .. |..... 4.. 8. 64. 128 ... ----------------------- -2|.1.. 2. 16. 32.... 0 .. |.... -2... 0 -32 ... ----------------- .... 1.. 0..16... 0 x^2 + 16 = 0 x^2 = -16 x = ±√ ̅(-16) = ±4√ ̅-1 This is not a real value. The equation is (x - 4)(x + 2)(x^2 + 16) = 0 Answers in real values are x = -2 x = 4

OpenStudy (anonymous):

IF this answer helped. Please click on the best response button :D